# What Matters to Enhance Traffic Rule Compliance of Imitation Learning for Automated Driving

### v2

@@ III-D Auxiliary Tasks
Auxiliary tasks have been proven to be efficient in many learning approaches. It has two main advantages. 1) By introducing auxiliary tasks, guarantees the important information flows remain in the networks. This information flow is critical for the decision network to make decisions. 2) Auxiliary tasks can guide the direction of gradient descent during training, which can lead the neural network to a more optimal location in the weight space. In this autonomous driving task, we would introduce two extra auxiliary tasks, namely traffic light classification and stop sign classification.
III-D1 Traffic Light Classification
@@
The output of the traffic light decoder should be a vector of 4 which indicates these four states red light, yellow light, green light, and none in the current frame. We then define $y_{\textrm{light}}$ as the ground truth traffic light vector of length 4 and $\hat{y}_{\textrm{light}}$ as the output of the traffic light decoder with the same shape. We use cross-entropy loss $\mathcal{L}_{\textrm{light}}=\mathcal{L}_{\textrm{CE}}(\hat{y}_{\textrm{light}},y_{\textrm{light}})$ for this task.
III-D2 Stop Sign Classification
@@
The output of the stop sign decoder should have a vector of 2 which indicates if a stop sign exists in the current frame. The ground truth stop sign vector of length 2 and the output of the stop sign decoder with the same shape are defined as $y_{\textrm{stop}}$ and $\hat{y}_{\textrm{stop}}$, respectively. Also, cross-entropy loss $\mathcal{L}_{\textrm{stop}}=\mathcal{L}_{\textrm{CE}}(\hat{y}_{\textrm{stop}},y_{\textrm{stop}})$ is used for this task
@@
III-E Penalty-based Imitation Learning
Our investigation revealed that the objective function used in imitation learning and the metric used for evaluating autonomous driving performance are not consistent with each other. Hence, achieving a low loss in the objective function does not necessarily guarantee a high driving score or successful route completion. We identified two possible factors contributing to this discrepancy.
@@ III-E Penalty-based Imitation Learning
  • The objective function is not sensitive to serious violations of the traffic rules, i.e. the violation of red lights and stop signs. The average objective function loss may not increase too much when violating the traffic rules despite that this violation may cause serious consequences which result in a huge drop in driving score and route completion.
Behavior Cloning (BC) is an imitation learning technique that seeks to replicate the behavior of an expert agent. As a result, the performance of the trained agent is inherently limited by the capabilities of the expert agent. In other words, if the expert agent makes a mistake, the trained agent will learn to replicate that mistake rather than correct it.
Our goal is to modify the objective function used in imitation learning to incorporate traffic rules. Specifically, we plan to introduce additional penalties (higher loss), when the agent generates short-term future waypoints that violate traffic rules during training.
Traffic rules can be expressed as constraint functions that define conditions the optimization problem must meet. In our approach, we focus on three specific aspects of driving behavior: running red lights, ignoring stop signs, and failing to slow down when turning. These are the primary issues we observed with our vanilla imitation learning method. To address these concerns, we propose three corresponding penalties that can be used to quantify and penalize these violations.
III-E1 Red Light Penalty
@@
${\mathcal{P}}_{\rm tl}=\mathbb{E}_{(\bm{s},\mathcal{W})\sim\mathcal{D}}[\mathbb{1}_{\rm red}\cdot\sum_{t=1}^{T}c_{i}\cdot{\rm max}\{0,y_{t}-\overline{d}\}]\text{,}$ (5)
where $y_{t}$ represents the $y$ axis value in the $t$-th predicted waypoints $\bm{\hat{w}_{t}}$ generated by the trained agent and $\overline{d}$ denotes the distance between the ego car and the stop line at the intersection along the forward direction of the ego car ($y$ axis). The weight parameter is denoted by $c_{i}$, and the sum of all weight parameters is equal to one. $\mathbb{1}_{\rm red}$ indicates the presence of a red light that may affect the agent in the current frame. $(\bm{s},\mathcal{W})$ represents the combination of observation state $\bm{s}$ and waypoints set $\mathcal{W}$ of one frame, while $\mathcal{D}$ refers to the entire dataset.
@@
III-E2 Stop Sign Penalty
@@
${\mathcal{P}}_{\rm ss}=\mathbb{E}_{(\bm{s},\mathcal{W})\sim\mathcal{D}}[\mathbb{1}_{\rm stopsign}\cdot{\rm max}\{v-\epsilon_{v},0\}]$ (6)
@@
The variables $\bm{\hat{w}_{0}}$ and $\bm{\hat{w}_{1}}$ represent the first and second predicted waypoints, respectively, while $\Delta t$ indicates the time interval between each frame during data collection. The function $\mathbb{1}_{\rm stopsign}$ serves as an indicator for stop sign checking. The maximum speed required to pass stop sign tests is denoted by $\epsilon_{v}$.
An upper-speed limit is established for an area affected by a stop sign, and only speeds lower than this limit are permitted for the agent to pass through. If the agent exceeds this speed limit, a penalty is imposed based on its speed. As the training network only generates predicted waypoints, the velocity estimated from the waypoints is used to compute the stop sign penalty.
III-E3 Speed Penalty
If the agent attempts to turn at excessive speed, a penalty will be enforced. The rationale behind this penalty is based on human driving experience, as it is commonly known that turning at high speeds can increase the risk of collisions with pedestrians or other objects due to reduced braking distance. The speed penalty is defined as follows:
${\mathcal{P}}_{\rm sp}=\mathbb{E}_{(\bm{s},\mathcal{W})\sim\mathcal{D}}[{\rm\sin}(\Delta\delta)\cdot{\max}\{v-v_{\textrm{lb}},0\}]$ (8)
@@ III-E3 Speed Penalty
III-E4 Objctive Function
We trained the network to minimize the difference between the predicted waypoints and the ground truth waypoints, applying an $l_{1}$-norm based loss function. Also, cross semantics generation and auxiliary tasks are trained together with the policy generation task. The final loss function is defined in Equation (4). By applying the penalties, we formalize the constrained optimization:
@@ III-E4 Objctive Function
$\min_{\bm{\Theta}}\quad$ $\mathcal{L}_{\textrm{final}}$ (9)
@@
This is the final objective function to optimize. For simplicity, these Lagrange multipliers $\lambda_{1}$, $\lambda_{2}$, $\lambda_{3}$ are considered fixed hyper-parameters. Well-chosen $\lambda_{1}$, $\lambda_{2}$, $\lambda_{3}$ are important for optimization. According to our experiments, too large hyper-parameters influence the behaviors in other scenarios while too small ones are not powerful enough for the agent to obey the corresponding traffic rules. The right part of Figure 1 demonstrates the process of penalty-based imitation learning in detail.
IV Experiments
IV-A Task Description
We focus on a navigation task where the vehicle navigates along predefined routes in various scenarios and areas. The vehicle is guided by GPS signals and situations where GPS signals are low or not available are not being considered. The predefined routes include specific scenarios to test the agent’s ability to respond to emergencies such as obstacle avoidance, other vehicles running red lights, and the sudden appearance of pedestrians on the road. The agent is required to complete the route within a specified time limit, and any time exceeding the limit is considered a failure in terms of route completion.
IV-B Training Dataset
Obtaining realistic driving data is a challenging task. Therefore, we opted to use the Carla simulator [22] to gather training data that has been processed by the expert policy. Our training dataset is the same as TransFuser [6], which comprises approximately 2500 routes through junctions in 8 different towns. These routes have an average length of 100 meters, and around 1000 routes along curved highways with an average length of 400 meters. We utilized the expert agent, similar to TransFuser, to generate this training data. The expert policy we use is the same export policy in TransFuser [6]. For each frame, We collect
@@ IV-B Training Dataset
  • The measurements of the steer, throttle, brake action values, and velocity in the current frame.
  • The traffic light information which includes the stop line position, and traffic light status.
@@
Note that the interval for each frame is 0.5 seconds, indicating 2FPS collection frequency.
IV-C Metrics
The metrics used to evaluate the behavior of each are Route completion and Infraction Scores.
IV-C1 Route Completion
Route completion (RC) refers to the proportion of the completed route out of the whole route. Suppose $R_{j}$ means the route completion proportion in route $j$ and $N_{\textrm{route}}$ means the number of routes. The RC can be defined by:
@@ IV-C1 Route Completion
IV-C2 Infraction Score
Infraction Score (IS) is used to measure the driving behavior of the agent. We define $p_{k}$ as the penalty for an infraction instance, $k$ is incurred by the agent and $n_{k}$ is the number of occurrences of infraction instance $k$. Then the Infraction Score can be defined by:
$IS=\prod_{k}^{\{\rm Ped,Veh,Stat,Red,Stop\}}s_{k}^{n_{k}}$ (12)
@@ IV-C2 Infraction Score
IV-C3 Driving Score
Driving Score (DS) aims to measure the overall driving performance. It is defined by the weighted average of the route completion with an infraction multiplier:
$DS=\frac{1}{N_{\textrm{route}}}\sum_{j}^{N_{\textrm{route}}}R_{j}IS_{j}$ (13)
where $R_{j}$ and $IS_{j}$ are the route completion and infraction score $IS$ (12) for $j$-th test route.
IV-C4 Infractions per km
Infractions per km refer to the average infraction numbers per kilometer. The infractions include collisions with pedestrians, vehicles, and static elements, running a red light, off-road infractions, route deviations, timeouts, and vehicle blocks.
${\rm Infractions\ per\ km}=\frac{\sum_{j}^{N_{\textrm{route}}}\#\ {\rm infractions}_{j}}{\sum_{j}^{N_{\textrm{route}}}l_{j}}$ (14)
where $l_{j}$ is the driven distance (in kilometers) for route $j$.
IV-D Test Results
We use Carla leaderboard - Town05 long benchmarks to evaluate our model. Town05 long benchmark contains 10 routes and all of these routes are over 2.5 km. This benchmark is also used by InterFuser [8] and TransFuser [21]. We also test our model in the Longest6 benchmark which is proposed in TransFuser+ [7].
Baseline. The other baselines we chose to compare with our model are TransFuser+, TransFuser, Geometric Fusion, and LateFusion. TransFuser [21] introduces the Transformer into the multi-sensor fusion architecture to achieve better end-to-end autonomous driving results. TransFuser+ [7], as an extension of TransFuser, leverages several auxiliary losses to ensure important information flows such as traffic light and road line information in the network. InterFuser [8] developed a safety control module to regulate the behaviors of the agent, preventing the agent from violating the traffic rules. LAV [9] utilizes data from both the ego vehicle and surrounding vehicles to enhance the dataset through data augmentation. This is achieved by learning a spatial intermediate representation that remains invariant to different viewpoints.
The evaluation results demonstrated in Table I indicate that our model has a huge increase in driving scores and infraction penalties compared to other SOTA models. As the table demonstrates, our model adheres to traffic rules more effectively in terms of collision, red light violations, and stop sign violations. Table II demonstrates the performance of our model in the longest6 benchmark provided by Transfuser+. Our model continues to surpass all other baselines in driving scores and infraction scores. Additionally, there has been a significant decrease in collisions with pedestrians and vehicles. While our model may not attain the top route completion score, its highest infraction score and driving score suggest that the model embraces a more cautious approach to mitigating traffic rule violations, ultimately leading to the highest driving score.
IV-E Ablation Studies
In this subsection, we analyze the influences of different penalty weights for corresponding traffic rules. As Table III demonstrates, two extra weights for each penalty are selected for comparison. We also provide the results of models without CSG and penalties for comprehensive analysis. We notice that the infractions of traffic lights and stop signs are largely reduced by adding penalties. Our proposed multi-sensor fusion technology (CSG) also decreases the possibility of hitting obstacles such as vehicles, pedestrians, and other statics. The results of different penalty weights are also listed in the table for comparison. We found that assigning greater weight to more severe violations will increase the performance of our model. For instance, we apply greater penalties for the red light and the stop sign violations compared to overspeeding by turning, since those two violations cause more serious consequences.
Model Driving score Route compl. Infrac. score Collision pedestrian Collision vehicle Collision static Red light infraction Stop Sign Infraction
@@ IV-E Ablation Studies
P-CSG(Ours) $\textbf{59.10}\pm 4.07$ $86.84\pm 2.99$ $\textbf{0.68}\pm 0.06$ $0.01\pm 0.01$ $\textbf{0.07}\pm 0.05$ $0\textbf{0.00}\pm 0$ 0.00 $0.04\pm 0.02$ $0\textbf{0.00}\pm 0$ 0.00
TransFuser [21] $34.50\pm 2.54$ $61.16\pm 4.75$ $0.56\pm 0.06$ $0.01\pm 0.01$ $0.58\pm 0.07$ $0.38\pm 0.05$ $0.12\pm 0.03$ $0.05\pm 0.02$
TransFuser+ [7] $36.19\pm 0.90$ $70.13\pm 6.80$ $0.51\pm 0.03$ $0\textbf{0.00}\pm 0$ 0.00 $0.40\pm 0.13$ $0.04\pm 0.03$ $0.11\pm 0.10$ $0.04\pm 0.01$
InterFuser [8] $50.64\pm 3.51$ $89.13\pm 4.12$ $0.57\pm 0.05$ $0\textbf{0.00}\pm 0$ 0.00 $0.09\pm 0.04$ $0.01\pm 0.01$ $\textbf{0.02}\pm 0.01$ $0.04\pm 0.01$
LAV [9] $45.20\pm 6.36$ $\textbf{91.55}\pm 5.61$ $0.49\pm 0.06$ $0\textbf{0.00}\pm 0$ 0.00 $0.92\pm 0.42$ $0.33\pm 0.50$ $0.28\pm 0.28$
TABLE I: Carla Leaderboard - Town 05 Long Benchmark. In our assessment of the Town05 long benchmark, we analyze several performance indicators including the Driving Score, Route Complication, Infraction Score, instances of collisions, violations of red light signals, and violations related to stop signs. These metrics are precisely outlined by the Carla leaderboard [22]. It’s important to highlight that all alternate baseline models undergo retraining and testing within our dedicated CARLA environment. Each model’s evaluation is conducted thrice, employing distinct random seeds on every occasion. The performance data of LAV is collected from their paper.
Model Driving score Route compl. Infrac. score Collision pedestrian Collision vehicle Collision static Red light infraction Stop Sign Infraction
@@
P-CSG(Ours) $\textbf{49.21}\pm 2.93$ $86.24\pm 2.46$ $\textbf{0.56}\pm 0.04$ 0.00 0.18 0.24 0.21 0.00
WOR [27] $20.53\pm 3.12$ $48.47\pm 3.86$ $0.56\pm 0.03$ 0.18 1.05 0.37 1.28 -
LAV [9] $32.74\pm 1.45$ $70.36\pm 3.14$ $0.51\pm 0.02$ 0.16 0.83 0.15 0.96 -
Late Fusion $22.47\pm 3.71$ $83.30\pm 3.04$ $0.27\pm 0.04$ 0.05 4.63 0.28 0.11 -
TransFuser+ [7] $47.30\pm 5.72$ $\textbf{93.38}\pm 1.20$ $0.50\pm 0.06$ 0.03 2.45 0.07 0.16 -
TABLE II: Carla Leaderboard - Longest6 Benchmark We assess our model with Longest6 benchmark provided by Transfuser+ [7]. This benchmark is also under the Carla leaderboard framework. The metrics of the driving score, route complication, infraction score, collisions, violations of red light signals, and violations related to stop signs are the same as the metrics in the Carla leaderboard. The performance data of all other baselines is collected in their paper.
Model Driving score Route compl. Infrac. score Collision pedestrian Collision vehicle Collision static Red light infraction Stop Sign Infraction

### v3

@@ 3.4 Auxiliary Tasks
Auxiliary tasks have been proven to be efficient in many learning approaches with two main advantages: 1) Guaranteeing the important information flows remain in the networks. This information flow is critical for the decision-making network. 2) Guiding the direction of gradient descent during training, leading the neural network to a more optimal location in the weight space. In this autonomous driving task, we introduce two extra auxiliary tasks: traffic light classification and stop sign classification.
3.4.1 Traffic Light Classification
@@
The output of the traffic light decoder should be a vector of 4, which indicates four states: red light, yellow light, green light, and none in the current frame. We then define $y_{\textrm{light}}$ as the ground truth traffic light vector of length 4 and $\hat{y}_{\textrm{light}}$ as the output of the traffic light decoder with the same shape. We use cross-entropy loss $\mathcal{L}_{\textrm{light}}=\mathcal{L}_{\textrm{CE}}(\hat{y}_{\textrm{light}},y_{\textrm{light}})$ for this task.
3.4.2 Stop Sign Classification
@@
The output of the stop sign decoder should have a vector of 1, indicating if a stop sign exists in the current frame. The ground truth stop sign vector of length 1 and the output of the stop sign decoder with the same shape are defined as $y_{\textrm{stop}}$ and $\hat{y}_{\textrm{stop}}$, respectively. Also, binary cross-entropy loss $\mathcal{L}_{\textrm{stop}}=\mathcal{L}_{\textrm{BCE}}(\hat{y}_{\textrm{stop}},y_{\textrm{stop}})$ is used for this task
@@
3.5 Penalty-based Imitation Learning with Constraint Optimization
Our investigation revealed that the objective function used in imitation learning and the metric used for evaluating autonomous driving performance are not consistent. Hence, achieving a low loss in the objective function does not necessarily guarantee a high driving score and route completion. We identified two possible factors contributing to this discrepancy.
@@ 3.5 Penalty-based Imitation Learning with Constraint Optimization
  • The objective function is not sensitive to serious violations of the traffic rules, i.e., the violation of red lights and stop signs. The average objective function loss may not increase too much when violating the traffic rules, although this violation may cause serious consequences, which result in a huge drop in driving score and route completion.
We aim to constrain the objective function of imitation learning to incorporate traffic rules. Traffic rules can be represented as constraint functions that specify the conditions the optimization problem must satisfy. Our approach focuses on three specific aspects: running red lights, ignoring stop signs, and failing to slow down when turning. These are the primary issues we observed with our vanilla imitation learning method. To address these issues, we propose three corresponding penalties that can be used to quantify and penalize these violations.
3.5.1 Red Light Penalty
@@
${\mathcal{P}}_{\rm tl}=\mathbb{1}_{\rm red}\cdot\sum_{t=1}^{T}c_{i}\cdot{\rm max}\{0,\hat{y}_{t}-y_{\textrm{stop}}\}\text{,}$ (5)
where $\hat{y}_{t}$ represents the $y$ axis value in the $t$-th predicted waypoints $\bm{\hat{w}_{t}}=(\hat{x}_{t},\hat{y}_{t})$ and $y_{\textrm{stop}}$ denotes the distance between the ego car and the stop line at the intersection along the forward direction of the ego car ($y$ axis). The weight parameter is denoted by $c_{i}$, and the sum of all weight parameters is equal to one. $\mathbb{1}_{\rm red}$ indicates the presence of a red light that may affect the agent in the current frame.
@@
3.5.2 Stop Sign Penalty
@@
${\mathcal{P}}_{\rm ss}=\mathbb{1}_{\rm stopsign}\cdot{\rm max}\{v-\epsilon_{v},0\}$ (6)
@@
The variables $\bm{\hat{w}_{0}}$ and $\bm{\hat{w}_{1}}$ represent the first and second predicted waypoints, respectively, while $\Delta t$ indicates the time interval between each frame. The function $\mathbb{1}_{\rm stopsign}$ serves as an indicator for stop sign checking. The maximum speed to pass stop sign tests is denoted by $\epsilon_{v}$.
An upper-speed limit $\epsilon_{v}$ (close to zero) is established for an area affected by a stop sign, and only speeds lower than this limit are permitted for the agent to pass through. If the agent exceeds this speed limit, a penalty is imposed based on its speed. As the training network only generates predicted waypoints, the speed estimated from the waypoints is used to compute the stop sign penalty.
3.5.3 Curvature Speed Penalty
A penalty will be enforced if the agent attempts to turn at excessive speed. The rationale behind this penalty is based on human driving experience, as it is commonly known that turning at high speeds can increase the risk of collisions with pedestrians or other objects due to longer braking distance. The speed penalty is defined as follows:
${\mathcal{P}}_{\rm sp}={\rm\sin}(\Delta\delta)\cdot{\max}\{v-v_{\textrm{lb}},0\}$ (8)
@@ 3.5.3 Curvature Speed Penalty
3.5.4 Objctive Function
We trained the network to minimize the difference between the predicted and ground truth waypoints, applying an $l_{1}$-norm-based loss function. Also, cross semantics generation and auxiliary tasks are trained together with the policy generation task. Equation (4) defines the final loss function. By applying the penalties, we formalize the constrained optimization:
@@ 3.5.4 Objctive Function
$\min_{\bm{\Theta}}$ $\mathcal{L}_{\textrm{final}}$ (9)
@@
This is the final objective function to optimize. For simplicity, these Lagrange multipliers $\lambda_{1}$, $\lambda_{2}$, $\lambda_{3}$ are considered fixed hyper-parameters. Well-chosen $\lambda_{1}$, $\lambda_{2}$, $\lambda_{3}$ are important for optimization. According to our experiments, hyperparameters that are too large influence behaviors in other scenarios while hyperparameters that are too small are not powerful enough for the agent to obey the corresponding traffic rules. The right part of Figure 2 demonstrates the process of penalty-based imitation learning in detail.
4 Experiments
4.1 Training Dataset
Obtaining realistic driving data is challenging. Therefore, we opted to use the CARLA simulator to gather training data that has been processed by the expert policy. Our training dataset comprises approximately 2500 routes through junctions in 8 different towns. These routes have an average length of 100 meters and around 1000 routes along curved highways with an average length of 400 meters. The expert policy we use is the updated one in TransFuser+. For each frame, We collect
@@ 4.1 Training Dataset
  • The steer, throttle, brake action values, and speed measurements in the current frame.
  • The traffic light information, which includes the stop line position and traffic light status.
@@
Note that the interval for each frame is 0.5 seconds, indicating 2FPS frequency.
4.2 Metrics
Route completion and infraction scores are used to evaluate the agent’s behavior.
Route Completion Route completion (RC) refers to the proportion of the completed route out of the whole route. Suppose $R_{j}$ means the route completion proportion in route $j$ and $N_{\textrm{route}}$ means the number of routes. The RC can be defined by:
@@ 4.2 Metrics
Infraction Score Infraction Score (IS) is used to measure the driving behavior of the agent. We define $p_{k}$ as the penalty for an infraction instance, $k$ is incurred by the agent, and $n_{k}$ is the number of occurrences of infraction instance $k$. Then, the infraction score can be defined by:
$IS=\prod_{k}^{\{\rm Ped,Veh,Stat,Red,Stop\}}s_{k}^{n_{k}}\text{.}$ (12)
@@
Driving Score Driving Score (DS) aims to measure the overall driving performance. It is defined by the weighted average of the route completion with an infraction multiplier:
$DS=\frac{1}{N_{\textrm{route}}}\sum_{j}^{N_{\textrm{route}}}R_{j}IS_{j}\text{,}$ (13)
where $R_{j}$ and $IS_{j}$ are route completion, infraction score $IS$ (12) for $j$-th route.
4.3 Test Results
We use CARLA Leaderboard - Town05 Long Benchmark and Longest6 Benchmark to evaluate our model. Town05 Long Benchmark contains 10 routes, all of which are over 2.5 km. This benchmark is also used by InterFuser, and TransFuser. Longest6 benchmark is proposed in TransFuser+ with increased traffic density, and challenging pre-crash traffic scenarios.
Model Driving score Route compl. Infrac. score
@@ 4.3 Test Results
P-CSG(Ours) $\textbf{59.10}\pm 4.07$ $86.84\pm 2.99$ $\textbf{0.68}\pm 0.06$
TransFuser [29] $34.50\pm 2.54$ $61.16\pm 4.75$ $0.56\pm 0.06$
TransFuser+ [12] $36.19\pm 0.90$ $70.13\pm 6.80$ $0.51\pm 0.03$
InterFuser [33] $50.64\pm 3.51$ $89.13\pm 4.12$ $0.57\pm 0.05$
LAV [6] $45.20\pm 6.36$ $\textbf{91.55}\pm 5.61$ $0.49\pm 0.06$
Table 1: CARLA Leaderboard - Town 05 Long Benchmark.
Model Driving score Route compl. Infrac. score
@@
P-CSG(Ours) $\textbf{49.21}\pm 2.93$ $86.24\pm 2.46$ $\textbf{0.56}\pm 0.04$
WOR [5] $20.53\pm 3.12$ $48.47\pm 3.86$ $0.56\pm 0.03$
LAV [6] $32.74\pm 1.45$ $70.36\pm 3.14$ $0.51\pm 0.02$
Late Fusion [35] $22.47\pm 3.71$ $83.30\pm 3.04$ $0.27\pm 0.04$
TransFuser+ [12] $47.30\pm 5.72$ $\textbf{93.38}\pm 1.20$ $0.50\pm 0.06$
Table 2: CARLA Leaderboard - Longest6 Benchmark.
The evaluation results in Table 2 indicate that our model significantly increases driving scores and infraction penalties compared to other baselines. Table 2 demonstrates the performance of our model in the Longest6 Benchmark provided by Transfuser+. Our model surpasses all other baselines in driving and infraction scores. While our model may not attain the top route completion score, its highest infraction score and driving score suggest that it embraces a more cautious approach to mitigating traffic rule violations, ultimately leading to the highest driving score.
4.4 Ablation Studies
In this subsection, we analyze the influences of different penalty weights for corresponding traffic rules. The ablation study focuses on significant hyper-parameters, namely $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$, which correspond to the penalty weights associated with the red light, speed, and stop sign. Our optimal model employs default weights of $\lambda_{1}=0.5$, $\lambda_{2}=0.05$, and $\lambda_{3}=0.5$. As Table 3 demonstrates,
Model Driving score Route compl. Infrac. score